Circular No. 8114

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

GRB 030329 AND SUPERNOVA 2003dh P. Garnavich, University of Notre Dame; T. Matheson, Harvard-Smith-

sonian Center for Astrophysics (CfA); E. W. Olszewski, University of Arizona; P. Harding, Case Western Reserve University; and K. Z. Stanek, CfA, report that additional spectra of the afterglow of GRB 030329 (IAUC 8101) and suspected supernova (IAUC 8108) were obtained with the 6.5-m MMT reflector on Apr. 8.13 UT. The spectral features reported by Garnavich et al. (ibid.) continue to develop. Subtracting a scaled version of the power-law spectrum of Apr. 4.27 from that of Apr. 8.13 reveals an energy distribution remarkably similar to that of SN 1998bw a week before maximum light (Patat et al. 2001, Ap.J. 555, 900). This is the first spectral confirmation linking supernovae to classical γ-ray bursts (GRBs). This spectrum can be

seen at http://cfa-www.harvard.edu/~tmatheson/compgrb.jpg.

R. Chornock, R. J. Foley, A. V. Filippenko, M. Papenkova, and D. Weisz, University of California, Berkeley, write: “Inspection of CCD spectra (range 310–1000 nm) of GRB 030329, obtained on Apr. 8 UT with the Shane 3-m telescope at Lick Observatory, confirms the emergence of broad bumps (especially at rest-frame 500 nm) characteristic of the peculiar type-Ic supernovae 1998bw and 2002ap at early times, as announced by Garnavich et al. on IAUC 8108. The association between core-collapse supernovae and at least some of the long-duration GRBs thus seems solid. We expect the supernova features to continue strengthening with time relative to the power-law continuum of the GRB afterglow.”

The designation 2003dh has been assigned to the supernova emerging from the afterglow of GRB 030329.

SUPERNOVA 2003di Further to IAUC 8112, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and K. Quinones report the discovery of an apparent supernova (mag 17.8) on unfiltered NEAT images taken on Feb. 4.58–4.60 UT. The new object is located at α = 12^{h}00^{m}24^{s}.08, δ = +11^{o}58′26′′.5 (equinox 2000.0), which is 17′′.1 east and 11′′.2 north of the center of the apparent host galaxy. Additional magnitudes for SN 2003di: 2002 Feb. 22.53, [20.6; Apr. 1.41, [20.5; 2003 Feb. 11.60, 17.1; 24.64, 17.2; Mar. 7.54, 17.9; 14.40, 18.0; 25.39, 18.4; Apr. 8.15, 19.2.

2003 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green